OpenStudy (anonymous):

lim(n→∞)⁡〖(n-1)/(n+3)〗= ? .. can't find the way to transform the equation be cause ∞/∞ is impossible...

OpenStudy (anonymous):

try writing down (n-1)/(n+3) as (1-1/n)/(1+3/n), i.e. dividing both numerator and denominator by n. That should help. :)

OpenStudy (anonymous):

ok thanks, first lesson at limits though..

OpenStudy (agreene):

\[\lim_{x\rightarrow\infty}\frac{n-1}{n+3}\] If that is actually the question--use l'Hopital's rule: where: \[\lim \frac{\infty}{\infty}=\lim_{x\rightarrow\infty}\frac{f'(x)}{g'(x)}\] So... \[\lim_{x\rightarrow\infty}\frac{\frac{d}{dn}(n-1)}{\frac{d}{dn}(n+3)}=\lim_{x\rightarrow\infty}\frac{1}{1}=1\]